Total Manufacturing Cost, Income Statement, Unit Cost, and Selling Price You are consulted by Investors, Inc., a group of investors planning a new product. They have estimates of the costs of materials, labor, overhead, and other expenses for 2016 but need to know how much to charge for each unit to earn a profit in 2016 equal to 10% of their estimated investment of $1,000,000 (ignore income taxes). Their plans indicate that each unit of the new product requires the following: Direct Material 4 lb. of a material costing $12 per lb. Direct labor 3hrs. of a die cutters time at $18 per hr 2 hrs. of an assemblers time at $16 per hr. Major tems of production overhead would be annual rent of s80,000 on the factory building and s50,000 on machinery as well as indirect material of $42.000. Other production overhead is an estimated 60% of total direct labor costs. Selling expenses are an estimated 20% of total sales, and non-factory administrative expenses are 10% of total sales. The consensus at Investors is that during 2016 4,000 units of product should be produced for selling and another 1,000 units should be produced for the next years beginning inventory. Also, an extra 6,000 pounds the manufacturing process, all units started must be completed, so work in process inventories are negligible of material will be purchased as beginning inventory for the next year. Because of the nature of Required
Menu a. Incorporate the above data into a schedule of estimated total manufacturing costs and compute the unit production cost for 2016. Do not use negative signs with any of your answers. Estimated Total Manufacturing Cost For the Year Ended December 31,2016 Direct material Beginning materials inventory Cost of material available Less: Total materials used Less: Direct materials used Direct labor
Indirect material Building rent Machinery rent Total manufacturing overhead Total manufacturing costs Round answer to two decimal places Product cost per unit 211.6
b. Prepare an estimated income statement that would provide the target amount of profit for 2016 Income Statement For the Year Ended December 31,2016 Sales Cost of goods sold Gross profit on sales Operating expenses: Selling expenses Administrative expenses Net income S100,000 c. What unit sales price should investors charge for the new product?
Answer #1

a. Schedule of estimated total manufacturing costs and unit production cost for 2016 is as below :

Cost Element Amount $ Amount $
Direct Material
Opening stock 0
Add, Purchase [(5000* 4lb + 6000lb)=26000lb @$12 312,000
Total available direct material 312,000
Less, Closing Stock ( 6000lbs @$12) 72,000
Direct Material issued to production 240,000
Direct Labour
Die cutter 5000units @ 3 hrs @$18 per hour 270,000
Assembler 5000units @ 2 hrs @$16 per hour 160,000 430,000
Variable production Overhead 60% of direct labour cost 258,000
Fixed Production overhead
Rent on factory building 80,000
Rent on machinery 50,000
Indirect material 42,000 172,000
Cost of Manufacturing 1,100,000
Unit Cost of production 1,100,000/5000 220
Add, Opening stock of finished good 0
Less, closing stock of finished good 1000 @ 220 220,000
Cost of goods sold 880,000

c. As unit sale price is not given, let us first answer question C first :

Let us assume unit sale price is = $ X

As such sales value will be = 4000 * $ X

Profit Required is 10% of $1,000,000 = 100,000

Selling and non-factory Admin expenses are

30% of sales = 0.3 * 4000 X = 1200X

Now the equation will be( in $) = 880,000 + 1200X + 100,000 = 4000X

hence 4000X - 1200X = 980,000

2800X = 980,000

hence X = 980,000/2800 = $350

As such unit sales price to be charged is $350

b. Estimated income statement that would provide the target amount of profit for 2016 is as below :

Description Amount $ Amount $ Amount $
Sales 4000 units @ $350 1,400,000
Less, Cost of Goods Sold
Direct Material
Opening stock 0
Add, Purchase [(5000* 4lb + 6000lb)=26000lb @$12 312,000
Total available 312,000
Less, Closing Stock ( 6000lbs @$12) 72,000
Direct Material issued to production 240,000
Direct Labour
Die cutter 5000units @ 3 hrs @$18 per hour 270,000
Assembler 5000units @ 2 hrs @$16 per hour 160,000 430,000
Variable production Overhead 60% of direct labour cost 258,000
Fixed Production overhead
Rent on factory building 80,000
Rent on machinery 50,000
Indirect material 42,000 172,000
Cost of Manufacturing 1,100,000
Unit Cost of production 1,100,000/5000 220
Add, Opening stock of finished good 0
Less, closing stock of finished good 1000 @ 220 220,000
Cost of goods sold 880,000
Gross Profit 520,000
Less, Selling and Non factory admin Expenses
Selling Expenses = 20% of sales 280,000
Admin Expenses = 10% of sales 140,000 420,000
Net Profit (Ignoring Income Tax) 100,000